Which solution has the lowest pOH? 

5.00x10-1 mol/L NH3

5.00x10-1 mol/L KOH

8.00x10-2 mol/L F-

8.00x10-2 mol/L Ca(OH)2 Chat started



1
Expert's answer
2021-03-03T06:29:30-0500

First, we need to write the chemical equation. HCN --> H+ + CN-




Next, make an ICE chart with the initial concentrations of [HCN] = 0.0620M and [H+] = [CN-] = 0. Assign a variable of x to the change in concentration for all three. This makes the equilibrium concentrations 0.0620-x, x, and x respectively.




Now, write an equilibrium expression, Ka = [H+][CN-]/[HCN]. Ka = 4.9E-10 = x2/(0.0620-x). Since the power of the Ka is -10, we can safely assume that the x is small enough to ignore compared to the 0.0620, so it becomes 4.9E-10=x2/0.0620. From that x (which is equilibrium [H+]), we can calculate pH = -log[H+].




Finally, watch out for the pH sig fig rule - the number of SFs in the concentration is the number of decimal places in the pH.